OpenStudy (hba):

An electron.starting from rest,has an acceleration that increases linearly with time;that is,a=kt,in which k=1.50 m/s^3 (a) Plot a versus t during during the first 10-s interval.(b)From the curve of part (a) plot the corresponding v versus t curve and estimate the electron's velocity 5 s after the motion starts.(c) From the v versus t curve of part (b) plot the corresponding x versus t curve and estimate how far the electron moves during the first 5 s of its motion.

OpenStudy (yrelhan4):

a vs t graph will be staright line passing through the origin with slope k.. its just like y=mx..

OpenStudy (hba):

What about the whole question?

OpenStudy (yrelhan4):

lol. wait. i just read the first part..

OpenStudy (yrelhan4):

for v... a=dv/dt right? so dv/dt=kt.. --> dv=kt*dt integrate both sides.. (delta)v = kt^2/2 . right?

OpenStudy (hba):

Dude this looks like a very long question to me so would you just start from the first part.

OpenStudy (yrelhan4):

haha you dont have to be rude man. i see you are an amby too :) i told you the first part part no? you get it would be a straight line passing through the origin? now what the question asking is.. fir 10 sec interval.. take a couple of values of t.. like 5 sec and 10 sec.. plot them.. and join them with the origin. thats it.. and.. you dont have think this is gonna be a long question. that slows down your brain..

OpenStudy (dls):

correct^

OpenStudy (hba):

@yrelhan4 I ain't being rude dude.So chillax. Okay i get it.

OpenStudy (yrelhan4):

I'm chillen, dont worry about it. :) And did you get what i did for finding v? and as it was an indefinite integral.. so no delta(v).. it would be v=1/2 kt^2 .. now plug in values of t. and find the corresponding values of v, and plot..

OpenStudy (hba):

Well yeah i am getting all of it.

OpenStudy (yrelhan4):

Thats good.. And for the last part.. write v=dx/dt.. integrate again.. plot the graph.. and the question is done.

OpenStudy (hba):

Thanks a ton.

OpenStudy (yrelhan4):

You're welcome.